The manager of a candy shop sells chocolate covered peanuts for $6 per pound and chocolate covered cashews for $9 per pound. The manager wishes to mix 60 pounds of the cashews to get a mixture that will sell for $8 per pound. How many pounds of peanuts should be used?
Answer by ankor@dixie-net.com(22740)

peanuts for $6 per pound and chocolate covered cashews for $9 per pound.
The manager wishes to mix 60 pounds of the cashews to get a mixture that will sell for $8 per pound.
How many pounds of peanuts should be used?
:
then
x = amt of peanuts required
:
A typical mixture equation
6x + 9(60) = 8(x+60)
6x + 540 = 8x + 480
540 - 480 = 8x - 6x
60 = 2x
x = 60/2
x = 30 lb of peanuts
:
:
Check solution
6(30) = 9(60) = 8(30+60)
180 + 540 = 8(90)